OpenStudy (anonymous):

0.7(3.6 x 10−2 ) in scientific notation

OpenStudy (anonymous):

What is the first thing you want to do? Do you know?

OpenStudy (anonymous):

1.4 (-1.+18. x)

OpenStudy (anonymous):

\[0.7×3.6×10^{−2}=2.52×10^{−2}\] @ohmygodhelp

OpenStudy (anonymous):

first off lets solve it .7(3.6*10^-2) exponents .7(3.6*.01) multiply .7(.036) multiply 0.0252 scientific notation 2.52*10^-2

OpenStudy (anonymous):

\[7x10^{-1}\left( 3.60 \times 1.00 × 10^{1} -2.00\right)\]